# V-Thinker: Interactive Thinking with Images

### v1

@@ 2 Related Work
Interactive visual reasoning is a long-term research vision. Early explorations, such as LLaVA-Plus [31] and Visual Sketchpad [24], pioneered this direction by enabling models to conduct visual operations during reasoning. These studies laid an important foundation for integrating interactive perception into multimodal reasoning systems. With the growing success of reinforcement learning (RL) in visual reasoning, OpenAI’s o3 [39] introduced the concept of thinking with images. Subsequent works [26, 62, 64, 83, 72, 6, 84, 71, 27, 80] have further developed this paradigm, provides a fresh perspective for the field of agentic RL. In particular, DeepEyes [81] and Thyme [79] employ RL-based training to guide models in performing executable visual tools (e.g., cropping) as part of their reasoning chain. Meanwhile, DeepSketcher [72] explores implicit visual reasoning, where models leverage abstract visual cues instead of explicit pixel-level manipulation. Building on this progress, our work advances interactive thinking with images by enabling the model to autonomously generate, execute, and iteratively refine visual code during reasoning. This opens a new perspective on integrating visual interaction into reasoning, paving the way toward more intuitive and human-like multimodal cognition.
@@ 5.2 Evaluation Dimensions
| Method | VTBench | General Reasoning | |||||||||||||
| Perception |
| GPT-4o | 2.3 | 3.7 | 38.3 | 43.8 | 68.8 | 26.3 | |||||||||
| InternVL3-78B | 10.8 | 16.0 | 43.4 | 43.1 | 64.2 | 27.7 | |||||||||
| InternVL3-8B | 10.4 | 5.4 | 36 | 29.3 | 58.8 | 24.9 | |||||||||
| LLaVA-OV-1.5-8B | 9.2 | 9.6 | 32.0 | 25.6 | 56.7 | 23.7 | |||||||||
| InternVL3-2B | 5.6 | 0.8 | 21.0 | 23.3 | 41.7 | 24.3 | |||||||||
[…]
@@ 6.1 Experimental Setup
For training, we conduct all experiments on H20 GPUs (i.e., 64 nodes with 8 GPUs per node). Both SFT stages use a learning rate of . The final SFT checkpoint initializes the RL stage, which is trained for one epoch with eight rollouts per iteration, a learning rate of , and a warm-up ratio of 0.05. For VTBench evaluation, we employ two models as judges: Qwen3-235B-A22B and Qwen3-VL-235B-A22B. Their results differ by less than 2% from GPT-4.1, validating their reliability as open-source evaluators under a unified evaluation protocol. For data construction, We employ GPT-5 as the generator and adopt Qwen3-VL-232B-A22B as the data checking module . Furthermore, GPT-4.1 is utilized as the repairer and perform progressive expansion.
@@ 6.2 Main Results
(2) Effectiveness of V-Thinker. Under identical experimental setups, V-Thinker consistently outperforms baseline LMMs across all three interactive reasoning domains, achieving an average accuracy improvement of 14.6% and maintaining over 8% gains on individual domains. Notably, it achieves over 25% performance improvement in the Instruction-Guided Interaction domain. These results underscore V-Thinker’s superior interactive thinking capability.
@@ 7 Conclusion
Limitation

### v2

@@ 2 Related Work
Interactive visual reasoning is a long-term research vision. Early explorations, such as LLaVA-Plus [31] and Visual Sketchpad [24], pioneered this direction by enabling models to conduct visual operations during reasoning. These studies laid an important foundation for integrating interactive perception into multimodal reasoning systems. With the growing success of reinforcement learning (RL) in visual reasoning, OpenAI’s o3 [39] introduced the concept of thinking with images. Subsequent works [26, 63, 65, 84, 73, 6, 85, 72, 27, 81, 47] have further developed this paradigm, provides a fresh perspective for the field of agentic RL. In particular, DeepEyes [82] and Thyme [80] employ RL-based training to guide models in performing executable visual tools (e.g., cropping) as part of their reasoning chain. Meanwhile, DeepSketcher [73] explores implicit visual reasoning, where models leverage abstract visual cues instead of explicit pixel-level manipulation. Building on this progress, our work advances interactive thinking with images by enabling the model to autonomously generate, execute, and iteratively refine visual code during reasoning. This opens a new perspective on integrating visual interaction into reasoning, paving the way toward more intuitive and human-like multimodal cognition.
@@ 5.2 Evaluation Dimensions
| Method | VTBench | General Reasoning | ||||||
| Perception | Instruct. Interaction | Interactive Reasoning | Avg. | MathVision Acc. | We-Math Acc. | VisuLogic Acc. | Avg. | |
| GPT-4o | 12.6 | 26.0 | 36.4 | 25.0 | 43.8 | 68.8 | 26.3 | 46.3 |
| InternVL3-78B | 13.8 | 19.0 | 43.4 | 25.4 | 43.1 | 64.2 | 27.7 | 45.0 |
| InternVL3-8B | 10.4 | 6.8 | 33.8 | 17.0 | 29.3 | 58.8 | 24.9 | 37.7 |
| LLaVA-OV-1.5-8B | 12.2 | 12.2 | 30.2 | 18.2 | 25.6 | 56.7 | 23.7 | 35.3 |
| InternVL3-2B | 3.0 | 3.4 | 22.0 | 9.5 | 23.3 | 41.7 | 24.3 | 29.8 |
| Qwen2.5-VL-7B | 12.6 | 8.8 | 31.8 | 17.7 | 23.0 | 61.7 | 26.0 | 36.9 |
| V-Thinker-7B | 18.6 | 31.6 | 40.4 | 30.2 | 29.3 | 62.8 | 26.6 | 39.6 |
| (vs Qwen2.5-VL-7B) | +6.0 | +22.8 | +8.6 | +12.5 | +6.3 | +1.1 | +0.6 | +2.7 |
@@ 6.1 Experimental Setup
For training, we conduct all experiments on H20 GPUs (i.e., 64 nodes with 8 GPUs per node). Both SFT stages use a learning rate of . The final SFT checkpoint initializes the RL stage, which is trained for one epoch with eight rollouts per iteration, a learning rate of , and a warm-up ratio of 0.05.
For VTBench evaluation, we employ two models as judges: Qwen3-235B-A22B and Qwen3-VL-235B-A22B. For Perception Tasks, the judges follow the prompts defined in Table 4, while for Instruction-Guided Interaction Tasks, they adopt the prompts specified in Table 5. Given the current limitations in model perception capabilities and the sensitivity of evaluation criteria to prompt variations, we fix the prompt version throughout all experiments to ensure consistent and fair judgments.
For data construction, We employ GPT-5 as the generator and adopt Qwen3-VL-232B-A22B as the data checking module . Furthermore, GPT-4.1 is utilized as the repairer and perform progressive expansion.
@@ 6.2 Main Results
(2) Effectiveness of V-Thinker. Under identical experimental setups, V-Thinker consistently outperforms baseline LMMs across all three interactive reasoning domains, achieving an average accuracy improvement of 12.5% and maintaining over 6% gains on individual domains. Notably, it achieves over 22% performance improvement in the Instruction-Guided Interaction domain. These results underscore V-Thinker’s superior interactive thinking capability.
@@ 7 Conclusion
Appendix
Contents
Appendix A Details of V-Thinker
A.1 Prompts Used in V-Thinker
Prompt Templates for Knowledge-Driven and Tool-Driven Synthesis.
To operationalize the two synthesis pathways in our co-evolution framework, we design two prompt specifications that explicitly encode how new instances are generated under knowledge-driven or tool-driven conditioning.
As shown in Table 6, the knowledge-driven prompt formalizes how sampled knowledge combinations guide the construction of problems, the rendering of figures, and the generation of reasoning trajectories involving visual tools. The underlying knowledge inventory is initialized from the 1,819 fundamental principles in We-Math 2.0 [43], and we instantiate both single-concept inputs and compositional combinations, so that the synthesized data can cover a broad range of reasoning objectives.
As shown in Table 7, the tool-driven prompt specifies the complementary mechanism in which visual tools act as the generative anchor. The initial visual tool system contains 61 visual tools, constructed by abstracting common operations from existing tool-augmented multimodal systems [26, 63, 65, 84, 73, 85, 27, 81, 80]. It defines how tool combinations constrain the image, the tool operations, and the resulting reasoning process, thereby enabling tool-conditioned data construction.
Together, these two prompt templates characterize how knowledge and tools drive generation along orthogonal dimensions. To ensure consistency across all visual tools, the rendering protocol enforces a unified set of conventions: all figures and tool overlays use absolute pixel coordinates, a common top-left origin, and full-bleed canvases without margins. These choices guarantee stable visual structure, precise localization, and reproducible execution across iterations. Further implementation details are provided in the corresponding prompt specifications.
Prompts for Coordinated Calibration
To support the coordinated calibration stage, we provide three checker prompt templates that separately evaluate (1) answer correctness, (2) integrity of the rendered image, and (3) consistency of auxiliary visual states. These templates are shown in Tables 8, 9, and 10.
For cases where only the textual answer is inconsistent while the visual components remain valid, a dedicated repair prompt (Table 11) reconstructs the question from the visual states to realign the reasoning. Together, these templates form the basis of the calibration mechanism used to refine the initial dataset into a coherent and verified collection.
Prompts for Progressive Expansion
Progressive expansion is operationalized through two coordinated prompt templates. The parallel extension prompt (Table 12) specifies how additional visual constructions are incorporated to introduce complementary observations beyond the original configuration. The sequential extension prompt (Table 13) prescribes constructions that are conditioned on existing entities or intermediate results, enabling reasoning to unfold through successive stages. These templates define the procedural rules for generating extended visual states and deeper reasoning trajectories, supporting the formation of higher-difficulty instances.
| Method | VTBench | |||
| Perception | Instruct. Interaction | Interactive Reasoning | Avg. | |
| GPT-4o | 12.6 | 26.0 | 36.4 | 25.0 |
| Qwen2.5-VL-72B | 38.0 | 34.2 | 51.4 | 41.2 |
| InternVL3-78B | 13.8 | 19.0 | 43.4 | 25.4 |
| InternVL3-8B | 10.4 | 6.8 | 33.8 | 17.0 |
| LLaVA-OV-1.5-8B | 12.2 | 12.2 | 30.2 | 18.2 |
| InternVL3-2B | 3.0 | 3.4 | 22.0 | 9.5 |
| Qwen2.5-VL-3B | 2.6 | 2.2 | 28.6 | 11.1 |
| LLaVA-OV-1.5-4B | 8.2 | 10.8 | 30.2 | 16.4 |
| Qwen2.5-VL-7B | 12.6 | 8.8 | 31.8 | 17.7 |
| V-Thinker-7B | 18.6 | 31.6 | 40.4 | 30.2 |
| (vs Qwen2.5-VL-7B) | +6.0 | +22.8 | +8.6 | +12.5 |
Prompts for Perception Data Synthesis
Perception-oriented data generation is supported by four prompt templates, corresponding to code construction and three levels of perceptual QA. The figure-construction prompt (Table 14) governs the generation of executable drawing code and the associated element-level annotations, ensuring that visual content is specified in absolute pixel coordinates and adheres to the structural constraints defined by element relations, element count, and sampled knowledge concepts.
Based on these annotations, three QA prompts convert visual structures into perception tasks along the hierarchical dimensions introduced above. The surface-level prompt (Table 15) produces localization questions grounded in explicit element coordinates. The semantic-level prompt (Table 16) builds questions around structurally defined feature points, requiring spatial interpretation beyond raw coordinates. The integrated reasoning prompt (Table 17) constructs tasks that combine perceptual identification with computational reasoning, such as determining derived geometric or structural quantities.
Together, these templates operationalize the construction of by transforming synthesized visual structures into systematically layered perceptual QA pairs.
Appendix B Additional Results and Analyses
B.1 Extended Quantitative Analysis
Table 3 provides an extended evaluation on VTBench by incorporating additional variants from the Qwen2.5-VL, InternVL3, and LLaVA-OneVision-1.5 families. Across this expanded comparison, two observations remain consistent.
Fine-grained perceptual grounding is consistently weak across model scales.
Large models such as GPT-4o, Qwen2.5-VL-72B, and InternVL3-78B exhibit limited accuracy on perception-oriented tasks, which reflects the difficulty posed by tasks that require local spatial grounding, such as identifying specific points, intersections, or geometric primitives.
V-Thinker shows consistent improvements in perception and interaction.
Relative to 7B-scale open-source models, V-Thinker achieves higher accuracy (+6.0%) on perception-oriented questions and a larger gain (+22.8%) on instruction-guided interaction, together with stable improvements in visual reasoning (+8.6%). These trends are consistently observed across the expanded baseline set.
B.2 Additional Analysis of the Data Evolution Flywheel
Expanded Knowledge Structures.
Figure 11 provides a comprehensive visualization of the knowledge system produced through the Data Evolution Flywheel. Beyond the summary shown in the main paper, the extended structure highlights the breadth and depth achieved through iterative evolution. The resulting hierarchy spans 25 domains and reaches up to 7 layers, forming a graph with 24,767 nodes. This expanded view further illustrates how the flywheel progressively enriches conceptual coverage, organizes related concepts into coherent clusters, and builds multi-level structures that support both fine-grained and cross-domain reasoning.
Tool System Analysis.
Within the Data Evolution Flywheel, the visual tool system co-evolves alongside the knowledge concepts, and the resulting tool set undergoes an additional consolidation stage. Since visual tools are instantiated through executable Python drawing routines, functionally identical operations may differ only in minor parameter settings (for example, line styles or rendering options), which can artificially inflate the apparent diversity of tools without reflecting meaningful distinctions in visual operations. To obtain a more faithful representation of the underlying tool space, we apply an additional round of BGE-based hierarchical clustering to the co-evolved tool set, using a strict similarity threshold of 0.3. The resulting clusters are then normalized through LMM-assisted (GPT-4.1) unification of tool names, followed by manual adjustment to ensure semantic consistency. This refinement phase collapses redundant variants and yields a compact, semantically coherent library of 234 visual tools, providing a clearer and more accurate characterization of the evolved visual tool system.
Appendix C Details of VTBench
C.1 Evaluation Dimensions
VTBench assesses interactive reasoning across three hierarchical dimensions, modeling the progression from perception to adaptive interaction:
Perception: Evaluates fine-grained visual perception, such as identifying or locating specific coordinates.
Instruction-Guided Interaction: Tests the ability to execute explicit visual instructions (e.g., drawing lines, labeling regions).
Interactive Reasoning: Evaluates reasoning tasks that require visual interaction.
C.2 Data Statistics and Evaluation Metrics
VTBench contains 1,500 QA pairs (500 per task type) across 9 open-source benchmarks (MathVista [34], MathVision [57], MathVerse [77], Dynamath [88], We-Math [40], LogicVista [67], CMM-Math [32], CharXiv [62] and ZeroBench [44]) covering four domains: Logical Reasoning, Geometry, Algebra, and Statistics. Evaluation is conducted under three task-specific metrics:
Perception Tasks: Models generate Python code to mark perceived coordinates. The resulting image is compared with annotations by large multimodal models (LMMs).
Instruction-Guided Interaction Tasks: Models generate Python code to perform instructed actions. The visual output is compared with expert annotations using LMM judgment.
Interactive Reasoning Tasks: Models output final answers, which are evaluated by LLMs for correctness.
Appendix D Detailed Experimental Setup
D.1 Implementation Details (Evaluation)
For VTBench evaluation, we employ two models as judges: Qwen3-235B-A22B and Qwen3-VL-235B-A22B. Their results differ by less than 2% from GPT-4.1, validating their reliability as open-source evaluators under a unified evaluation protocol. For all other benchmarks used in this work, including MathVision, VisuLogic, and We-Math, we follow the official evaluation procedures specified by each benchmark, including the provided scoring scripts and answer-matching rules.
D.2 Details of Baselines.
We evaluate all models on MathVision, VisuLogic, We-Math and VTBench, following each benchmark’s official protocol and reporting accuracy as the primary metric. Our baseline comparison includes a broad spectrum of multimodal reasoning systems. GPT-4o is OpenAI’s flagship multimodal model designed for unified vision–language understanding with strong cross-domain reasoning capability [38]. The Qwen2.5-VL family (72B/7B/3B) is an open-source vision–language series that emphasizes visual–language alignment and multimodal reasoning [3]. The InternVL3 series (78B/8B/2B) represents an open-source vision–language architecture with enhanced multimodal perception and reasoning capability [86]. The LLaVA-OneVision-1.5 family (8B/4B) consists of open multimodal models incorporating a unified vision encoder and language backbone, optimized for efficient and competitive vision–language reasoning [2]. Together, these baselines span a wide range of architectures and parameter scales, enabling a comprehensive and balanced comparison against our model.
Appendix E Broaden Impact
Advancing interactive, vision-centered reasoning in multimodal systems.
Our work encourages a transition from passive visual perception to interactive reasoning within images. Instead of only understanding visual content, V-Thinker can actively engage with visual elements through pointing, annotating, and manipulating structured regions. This shift supports explicit intermediate reasoning steps and improves transparency in visual-language decision making. Such interactive capabilities may influence a wide range of tasks including diagram understanding, scientific figure analysis, embodied perception, and human–AI collaborative interfaces, where traceable and verifiable reasoning is essential.
Enabling a creator-oriented synthetic data paradigm that expands future possibilities.
V-Thinker revisits the long-standing paradigm in which models act purely as solvers when synthesizing data. This solver-centric view restricts the diversity and structural richness of generated samples, particularly for tasks requiring precise spatial or logical alignment. Our work reveals that modern multimodal models can instead function as creators capable of generating complex visual problems, programmatically rendered images, auxiliary diagrams, and coherent reasoning paths. When integrated with knowledge-driven representations, this creator-oriented paradigm significantly enlarges the design space for synthetic reasoning data. It reduces reliance on handcrafted seeds, enables scalable and controlled curriculum evolution, and offers a foundation for autonomous dataset construction and simulation-based training. This paradigm shift may inspire future research on model-driven data ecosystems, controllable synthetic corpora, and new forms of interaction-centric supervision for robust multimodal intelligence.
Bridging foundational models with practical, tool-oriented applications.
The structured visual interactions and interpretable reasoning processes supported by V-Thinker help narrow the gap between multimodal foundation models and their real-world deployment. The ability to generate, manipulate, and reason with structured visual information can benefit scientific analysis, education technologies, robotics perception, and interactive decision-support systems. The explicit reasoning traces also promote safer and more accountable AI behavior, which is crucial in settings that demand correctness and transparency. By supporting precise visual alignment and interpretable reasoning, our work contributes to building reliable multimodal systems that operate effectively in complex visual environments.
Appendix F Limitation